A running 61-kg cheetah has a top speed of 28 m/s. (a) What is the cheetah's maximum kinetic energy? J (b) Find the cheetah's speed when its kinetic energy is one half of the value found in part (a). m/s
Answer #1

The maximum kinetic energy of the cheetah is, KEmax (61.0kg) (28.0m/s) 2.398 104 J The speed of the cheetah when it is having
Answer #2

SOLUTION :


a.


Maximum K.E. of cheetah 

= 1/2 * 61 * 28^2 

= 23912 J (ANSWER).


b.


If K.E. of cheetah is half of that in (a) above,


New K. E. of cheetah = 1/2 * 61 * V^2

=> 23912 / 2 = 30.5 V^2

=> V = sqrt(23912/61)

=> V = new speed of cheetah = 19.8 m/s (ANSWER).

answered by: Tulsiram Garg